OpenStudy (anonymous):

1+3=2 how is this possible?

OpenStudy (anonymous):

i don't think its possible

OpenStudy (anonymous):

it could be a -1 + 3

OpenStudy (anonymous):

nope it 1+3=2 and it is possible

OpenStudy (anonymous):

no variables you might have missed?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i have no answer than:o

OpenStudy (anonymous):

1 + 3 = 2 move 2 to the left by subtracting -2 -2 1 + 1 add so 2? answer i dont think this could be it though

OpenStudy (anonymous):

nope not quite but close

OpenStudy (anonymous):

here's a hint ignore the 2 and mess around with the equation to get 2 different answers that when added together you get 2

OpenStudy (anonymous):

i give up:/

OpenStudy (anonymous):

1+3=4 1-3=-2 4-2=2

OpenStudy (anonymous):

how did you get -3 and -2?

OpenStudy (kainui):

Lol this is ridiculous. Go back to elementary school @mathman806

OpenStudy (anonymous):

it's a riddle

OpenStudy (cggurumanjunath):

well done it was really cool. @mathman806 .

OpenStudy (kainui):

No it was poorly done, the answer was nonsense.

OpenStudy (anonymous):

yes it was poorly done, sadly

OpenStudy (cggurumanjunath):

it was just a riddle.

OpenStudy (kainui):

Where's the riddle? Explain the riddle @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

chill dude @kainui

OpenStudy (kainui):

Good explanation. =) @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

thx.lol. @Kainui